OpenStudy (anonymous):

Martina combined different amounts of water, vinegar, and rubbing alcohol to make 38 milliliters (ml) of a solution for a science experiment. The amount of water was x ml. The amount of vinegar was 8 ml less than that of water, and the amount of rubbing alcohol was half the amount of vinegar. This is represented in the equation below: x + (x − 8) + one half (x − 8) = 38 What was the difference in the amounts of water and rubbing alcohol Martina combined? 8 ml 14 ml 18 ml 20 ml

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

x + x - 8 + x/2 - 4 = 38 2x + x/2 = 38 + 8 + 4 5x/2 = 50 5x = 100 x = 20 so d?

OpenStudy (anonymous):

That's the amount of water: 20 The amount of rubbing alcohol: (1/2)(20-8) = (1/2)(12) = 6 Difference: 20-6 = 14 not d but B

imqwerty (imqwerty):

yes correct :)

OpenStudy (anonymous):

:D ok ty for checkin

imqwerty (imqwerty):

np (: